OpenStudy (anonymous):

Jeff is designing a seating arrangement for a concert. To give everyone a good view each row must have 5 more seats than the row before it, and the first row can only have 12 seats. Help Jeff plan the rest of the seating by solving for how many seats are in row 25. Then explain how to create an equation to predict the # of seats in any row.

OpenStudy (anonymous):

@thanooz would you mind helping me? I'll give a medal

OpenStudy (anonymous):

now let no of seats in first row be 'x' now no of seats in second row ,i.e, in next row be ""x+5"" ="" x+(1)5 "" =""x+(2-1)5"" now no of seats in third row are"""x+5 +5"""="x+10" ='"x+(2)5"' ="'x+(3-1)5"' similarly, no of seats in nth row are N=''x+(n-1)5'' where N= no of seats............

OpenStudy (anonymous):

Thus no of seats of 25th row are N=x+(n-1)5 =12+(25-1)5 =12+(24)5 =12+120 there fore N =132